# dark matter and baryon asymmetry of the Universe

### v1

@@ I Introduction
In Ref. Aoki:2022bkg , an extension of the AKS model was examined. It has a general Yukawa interaction between the SM fermions and the Higgs doublets, while those in the original AKS model are constrained by a softly broken $Z_{2}$ symmetry to avoid the flavor-changing neutral current (FCNC). By relaxing this symmetry, the model includes multiple $CP$-violating phases in not only the Higgs potential but also the Yukawa interactions, which make it possible to avoid the current severe constraint on the electron electric dipole moment (eEDM) ACME:2018yjb ; Roussy:2022cmp by the destructive interference between these $CP$-violating phases Kanemura:2020ibp . In Ref. Aoki:2022bkg , the authors performed a numerical analysis of the baryon number production via EWBG in the model and found one benchmark scenario where tiny neutrino masses, DM and the BAU can be simultaneously explained while avoiding all current experimental and theoretical constraints. However, Ref. Aoki:2022bkg assumes ad hoc flavor-alignment structure in the Yukawa interaction to restrict the FCNC Pich:2009sp , while the original AKS model naturally prohibits it at tree level by the softly broken $Z_{2}$ symmetry Glashow:1976nt . In addition, relaxing this symmetry causes too many model parameters, which makes it difficult to give a solid prediction.
@@
In this paper, according to the facts mentioned above, we revisit the original AKS model to investigate the $CP$ violation. We examine the neutrino mass, lepton-flavor-violating (LFV) processes, DM, the eEDM, and other experimental and theoretical constraints. We evaluate the eEDM induced by a new Yukawa interaction between the charged leptons and the dark sector particles, which was neglected in Ref. Aoki:2022bkg . We consider a scenario where the eEDM is suppressed by the destructive interference between the dark sector contribution and the Higgs sector contribution and avoids the current severe constraint. This is a different suppression mechanism from that employed in Ref. Aoki:2022bkg , which was proposed in Ref. Kanemura:2020ibp and studied in Ref. Kanemura:2021atq ; Enomoto:2021dkl ; Kanemura:2023juv ; Kanemura:2023jbz . It is also different from other conventional suppression mechanisms in the two Higgs doublet models Inoue:2014nva ; Modak:2018csw ; Fuyuto:2019svr ; Bian:2014zka ; Bian:2016zba ; Abe:2013qla ; Cheung:2020ugr ; Altmannshofer:2020shb .
@@
We show two benchmark scenarios where tiny neutrino masses and DM can be explained simultaneously while avoiding all current experimental and theoretical constraints. In the two scenarios, DM is the real scalar boson, but with different masses. In the first scenario, the mass is about half of the 125 GeV Higgs boson, and a pair of DM mainly annihilates into the SM fermions via an almost on-shell Higgs boson. On the other hand, in the second scenario, DM has heavier mass, and a pair of them can annihilate into the additional Higgs bosons. In both scenarios, the size of the eEDM is smaller than the current limit while keeping a large $CP$-violating phase ($\sim\mathcal{O}(1)$) in the Higgs potential, which is necessary for successful EWBG. Although our ultimate goal is to build a scenario of the model where the BAU can also be explained in addition to the neutrino mass and DM, here, we do not discuss EWBG in detail to avoid making the paper lengthy. The evaluation of the baryon number density produced by EWBG will be performed in our following paper.
@@ II.1 Lagrangian
$V_{\mathrm{THDM}}$ is the potential in the $CP$-violating two Higgs doublet model;
@@
which is the same as in the Type-X two Higgs doublet model Logan:2009uf ; Su:2009fz ; Grossman:1994jb ; Aoki:2009ha ; Barger:1989fj . After the bi-unitary transformations of the SM fermions, the Yukawa interactions are given by
@@
We note that the coupling $h_{i}^{\alpha}$ is defined for the mass eigenstates of the charged leptons (without ${}^{\prime}$). The matrix $h$ includes nine complex parameres. Three of them can be real by rephasing the lepton fields. This degree of freedom is the same as those to reduce the three unphysical phases in the (Pontecorvo-Maki-Nakagawa-Sakata) matrix $U$ ref: Pontecorvo ; Maki:1962mu . We have a detailed discussion about in Sec. III.1.
@@ II.2 Theoretical constraints
In Eq. (28), we use $|\lambda_{5}|$, not $\lambda_{5}^{R}$ shown in Ref. Aoki:2011zg , to include the effect of the $CP$-violation in $\lambda_{5}$ according to discussion in the two Higgs doublet models Deshpande:1977rw ; Klimenko:1984qx ; Nie:1998yn ; Ferreira:2004yd .
@@
The above is a rough evaluation of the theoretical constraint on the scalar potential in the model. More detailed analysis is necessary for discussing the non-decoupling effect of the additional scalar bosons Kanemura:2002vm ; Kanemura:2004mg ; Braathen:2019pxr , which causes the strongly first-order electroweak phase transition Kanemura:2004ch ; Kakizaki:2015wua ; Hashino:2016rvx ; Hashino:2021qoq ; Hashino:2022tcs ; Dorsch:2016nrg ; Enomoto:2021dkl ; Kanemura:2023juv ; Cline:2011mm necessary for successful electroweak baryogenesis Dorsch:2016nrg ; Enomoto:2021dkl ; Kanemura:2023juv ; Cline:2011mm ; Aoki:2022bkg . It will be performed in our next paper.
@@ III.1 Neutrino mass
$O=$ $\frac{1}{c_{\psi_{1}}c_{\psi_{2}}c_{\psi_{3}}}\begin{pmatrix}1&0&0\\ 0&c_{\gamma_{1}}&s_{\gamma_{1}}\\ 0&-s_{\gamma_{1}}&c_{\gamma_{1}}\\ \end{pmatrix}\begin{pmatrix}c_{\psi_{1}}&0&0\\ 0&1&is_{\psi_{1}}\\ 0&-is_{\psi_{1}}&1\\ \end{pmatrix}\begin{pmatrix}c_{\gamma_{2}}&0&s_{\gamma_{2}}\\ 0&1&0\\ -s_{\gamma_{2}}&0&c_{\gamma_{2}}\\ \end{pmatrix}$ (61)
@@ III.2 Lepton flavor violation
$(\sigma v)_{\gamma}=$ $\frac{\alpha^{2}}{256\pi^{3}s}\int_{-1}^{1}\mathrm{d}\cos\theta\Bigl{\{}s^{2}|A_{\gamma\gamma}|^{2}+2s^{2}m_{\eta}^{2}\mathrm{Re}[A_{\gamma\gamma}B_{\gamma\gamma}^{\ast}]+|B_{\gamma\gamma}|^{2}\Bigl{(}(m_{\eta}^{2}-ut)+s^{2}m_{\eta}^{4}\Bigr{)}\Bigr{\}},$ (86)
@@ IV Electric dipole moments
Next, we discuss the contribution from the Higgs sector $d_{e}^{\theta_{5}}$. This is the same with the eEDM in the THDM studied in Refs. Inoue:2014nva ; Modak:2018csw ; Fuyuto:2019svr ; Bian:2014zka ; Bian:2016zba ; Abe:2013qla ; Cheung:2020ugr ; Altmannshofer:2020shb . It is known that one-loop contribution is tiny because of the mass suppression by $m_{e}$, and the leading contribution comes from Barr-Zee type diagrams at two-loop level Barr:1990vd . We employ the formula in Ref. Altmannshofer:2020shb , which includes all the two-loop diagrams and gives the gauge-invariant result.
@@ V.4.1 Scenario I
$R\simeq\begin{pmatrix}0.997&0.0657&-0.0307\\ -0.0433&0.880&0.474\\ 0.0581&-0.471&0.880\\ \end{pmatrix}.$ (118)
@@
The original scalar quartic couplings in $V_{THDM}$ is evaluated by using the above input values as follows;
@@
We have checked that these quartic couplings satisfy the conditions for the vacuum stability in Eqs. (27)–(28) and those for the perturbative unitarity in Eqs. (32)–(37). In addition, we expect that they are small enough to satisfy the triviality bound up to $10~{}\mathrm{TeV}$ Kanemura:2023wap ; Dorsch:2016nrg .
@@ V.4.2 Scenario II
The other input parameters for $V_{THDM}$ are the same as in Scenario I. This change of $\theta_{5}$ leads to a slightly different mass spectrum;
@@ VI Discussions and conclusion
In this paper, we have not evaluated the EWPT and EWBG in the model. Here, we shortly comment about them although a detailed investigation will be performed in our following paper. For successful electroweak baryogenesis, the EWPT needs to be strongly first-order Kuzmin:1985mm . To achieve it in the model, the non-decoupling effect of the additional scalar bosons is required Kanemura:2004ch ; Kakizaki:2015wua ; Hashino:2016rvx ; Hashino:2021qoq ; Hashino:2022tcs ; Dorsch:2016nrg ; Enomoto:2021dkl ; Kanemura:2023juv ; Cline:2011mm ; Aoki:2022bkg . The non-decoupling effect also induces the sizable deviation in the Higgs triple coupling Kanemura:2002vm ; Kanemura:2004mg ; Braathen:2019pxr , which can be tested in the process of di-Higgs production at future high-energy colliders Cepeda:2019klc ; Fujii:2015jha ; Bambade:2019fyw . In our benchmark scenarios, by taking $\mu_{S}=20~{}\mathrm{GeV}$, we have $R_{hhh}\equiv\Delta\lambda_{hhh}/\lambda_{hhh}\simeq 22~{}\%$, which would be tested at the upgraded ILC with the beam energy $\sqrt{s}=1~{}\mathrm{TeV}$, where the expected sensitivity is $10~{}\%$. We also have checked that the strongly first-order EWPT can be realized in this case by using the one-loop effective potential Coleman:1973jx ; Dolan:1973qd . In addition, since $\theta_{5}$ is large ($\theta_{5}\sim\mathcal{O}(1)$) in both benchmark scenarios, we expect that the EWBG scenario can work. We will perform a more precise investigation of the estimation of the BAU and the theoretical constraints in our next paper.
@@ Acknowledgements
The work of K. E. was supported in part by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) No. NRF-2021R1A2C2009718. The work of S. K. was supported in part by the JSPS KAKENHI Grant No. 20H00160 and No. 23K17691.

### v2

@@ I Introduction
In Ref. Aoki:2022bkg , an extension of the AKS model was examined. It has a general Yukawa interaction between the SM fermions and the Higgs doublets, while those in the original AKS model are constrained by a softly broken $Z_{2}$ symmetry to avoid the flavor-changing neutral current (FCNC). By relaxing this symmetry, the model includes multiple $CP$-violating phases in not only the Higgs potential but also the Yukawa interactions, which make it possible to avoid the current severe constraint on the electron electric dipole moment (eEDM) ACME:2018yjb ; Roussy:2022cmp by the destructive interference between these $CP$-violating phases Kanemura:2020ibp . In Ref. Aoki:2022bkg , the authors performed a numerical analysis of the baryon number production via EWBG in the model and found one benchmark scenario where tiny neutrino masses, DM and the BAU can be simultaneously explained while avoiding all current experimental and theoretical constraints. However, Ref. Aoki:2022bkg assumes ad hoc flavor-alignment structure in the Yukawa interaction to restrict the FCNC Pich:2009sp , while the original AKS model naturally prohibits it at tree level by the softly broken $Z_{2}$ symmetry Glashow:1976nt . In addition, relaxing this symmetry causes too many model parameters.
@@
In this paper, according to the facts mentioned above, we revisit the original AKS model to investigate the $CP$ violation. The most significant difficulty is avoiding the constraint from the eEDM measurements unless we assume that $CP$-violating phases are very small. We investigate the eEDM in the AKS model to see if it is possible to avoid the current severe bound while keeping a large $\mathcal{O}(1)$ $CP$-violating phase in the Higgs potential, which is necessary for successful EWBG Turok:1990zg ; Cline:1995dg ; Fromme:2006cm ; Cline:2011mm ; Enomoto:2021dkl ; Kanemura:2023juv ; Aoki:2022bkg ; Dorsch:2016nrg . The eEDM is induced by two kinds of Feynman diagrams at the leading order. One is generated by the additional Higgs bosons like in the two Higgs doublet models (THDMs) Inoue:2014nva ; Modak:2018csw ; Fuyuto:2019svr ; Bian:2014zka ; Bian:2016zba ; Abe:2013qla ; Cheung:2020ugr ; Altmannshofer:2020shb , which depends on a $CP$-violating phase in the Higgs potential. The other one is generated by a new $CP$-violating Yukawa interaction between the charged leptons and the dark sector particles, which was neglected in Ref. Aoki:2022bkg . We consider a scenario where the eEDM is suppressed by the destructive interference between the Higgs sector contribution and the dark sector contribution to avoid the current severe constraint. This is a different suppression mechanism from that employed in Ref. Aoki:2022bkg , which was proposed in Ref. Kanemura:2020ibp and has been studied in Ref. Kanemura:2021atq ; Enomoto:2021dkl ; Kanemura:2023juv ; Kanemura:2023jbz . It is also different from other conventional suppression mechanisms in the THDMs Inoue:2014nva ; Modak:2018csw ; Fuyuto:2019svr ; Bian:2014zka ; Bian:2016zba ; Abe:2013qla ; Cheung:2020ugr ; Altmannshofer:2020shb .
@@
We also examine the neutrino mass, lepton-flavor-violating (LFV) processes, and DM in the model and show two benchmark scenarios where tiny neutrino masses and DM can be simultaneously explained while avoiding all current experimental and theoretical constraints. In the two scenarios, DM is the real scalar boson, but with different masses. In the first scenario, the mass is about half of the 125 GeV Higgs boson, and a pair of DM mainly annihilates into the SM fermions via an almost on-shell Higgs boson. On the other hand, in the second scenario, DM has heavier mass, and a pair of them can annihilate into the additional Higgs bosons. In both scenarios, the size of the eEDM is smaller than the current limit while keeping an $\mathcal{O}(1)$ $CP$-violating phase in the Higgs potential.
@@ II.1 Lagrangian
$V_{\mathrm{THDM}}$ is the potential in the $CP$-violating THDM;
@@
which is the same as in the Type-X THDM Logan:2009uf ; Su:2009fz ; Grossman:1994jb ; Aoki:2009ha ; Barger:1989fj . After the bi-unitary transformations of the SM fermions, the Yukawa interactions are given by
@@
We note that the coupling $h_{i}^{\alpha}$ is defined for the mass eigenstates of the charged leptons (without ). The matrix $h$ includes nine complex parameres. Three of them can be real by rephasing the lepton fields. This degree of freedom is the same as those to reduce the three unphysical phases in the (Pontecorvo-Maki-Nakagawa-Sakata) matrix $U$ ref: Pontecorvo ; Maki:1962mu . We have a detailed discussion about in Sec. III.1.
@@ II.2 Theoretical constraints
In Eq. (28), we use $|\lambda_{5}|$, not $\lambda_{5}^{R}$ shown in Ref. Aoki:2011zg , to include the effect of the $CP$-violation in $\lambda_{5}$ according to discussion in the THDMs Deshpande:1977rw ; Klimenko:1984qx ; Nie:1998yn ; Ferreira:2004yd .
@@ III.1 Neutrino mass
$O=$ $\frac{1}{c_{\psi_{1}}c_{\psi_{2}}c_{\psi_{3}}}\begin{pmatrix}1&0&0\\ 0&c_{\gamma_{1}}&s_{\gamma_{1}}\\ 0&-s_{\gamma_{1}}&c_{\gamma_{1}}\\ \end{pmatrix}\begin{pmatrix}c_{\psi_{1}}&0&0\\ 0&1&is_{\psi_{1}}\\ 0&-is_{\psi_{1}}&1\\ \end{pmatrix}\begin{pmatrix}c_{\gamma_{2}}&0&s_{\gamma_{2}}\\ 0&1&0\\ -s_{\gamma_{2}}&0&c_{\gamma_{2}}\\ \end{pmatrix}$
@@ III.2 Lepton flavor violation
$(\sigma v)_{\gamma}=$ $\frac{\alpha^{2}}{256\pi^{3}s}\int_{-1}^{1}\mathrm{d}\cos\theta\Bigl{\{}s^{2}|A_{\gamma\gamma}|^{2}+2s^{2}m_{\eta}^{2}\mathrm{Re}[A_{\gamma\gamma}B_{\gamma\gamma}^{\ast}]+|B_{\gamma\gamma}|^{2}\Bigl{(}(m_{\eta}^{4}-ut)^{2}+s^{2}m_{\eta}^{4}\Bigr{)}\Bigr{\}},$ (67)
@@ IV Electric dipole moments
Next, we discuss the contribution from the Higgs sector $d_{e}^{\theta_{5}}$. This is the same with the eEDM in the THDMs studied in Refs. Inoue:2014nva ; Modak:2018csw ; Fuyuto:2019svr ; Bian:2014zka ; Bian:2016zba ; Abe:2013qla ; Cheung:2020ugr ; Altmannshofer:2020shb . It is known that one-loop contribution is tiny because of the mass suppression by $m_{e}$, and the leading contribution comes from Barr-Zee type diagrams at two-loop level Barr:1990vd . We employ the formula in Ref. Altmannshofer:2020shb , which includes all the two-loop diagrams and gives the gauge-invariant result.
@@ V.4.1 Scenario I
$R\simeq\begin{pmatrix}0.997&0.0657&-0.0307\\ -0.0433&0.880&0.474\\ 0.0581&-0.471&0.880\\ \end{pmatrix}.$
@@
The original scalar quartic couplings in $V_{\mathrm{THDM}}$ is evaluated by using the above input values as follows;
@@
We have checked that these quartic couplings satisfy the conditions for the vacuum stability in Eqs. (27)–(28) and those for the perturbative unitarity in Eqs. (32)–(37). In addition, we expect that they are small enough to satisfy the triviality bound up to $10~{}\mathrm{TeV}$ Kanemura:2023wap .
@@ V.4.2 Scenario II
The other input parameters for $V_{\mathrm{THDM}}$ are the same as in Scenario I. This change of $\theta_{5}$ leads to a slightly different mass spectrum;
@@ VI Discussions and conclusion
In this paper, we have investigated two kinds of contributions to the eEDM at the leading order. The size of the next leading order (NLO) contribution is expected to be suppressed by $\mathcal{O}(\alpha)$. In the benchmark scenarios, it is roughly estimated by $d_{e}^{\mathrm{NLO}}=\mathcal{O}(10^{-31})$, which is one order smaller than the current experimental bound. Thus, the interference between $d_{e}^{h}$ and $d_{e}^{\theta_{5}}$ is not largely changed by the NLO contribution. We expect that $\theta_{5}$ can be $\mathcal{O}(1)$ even considering the eEDM at the NLO.
@@
Finally, we shortly comment on EWBG in the model. For successful EWBG, we need to satisfy the second and third conditions of the Sakharov condition by new physics, i.e., large enough $CP$ violation and departure from thermal equilibrium Sakharov:1967dj . $\theta_{5}$ is a source of $CP$ violation in the model. It is $\mathcal{O}(1)$ in the benchmark scenarios and is expected to be large enough to generate the observed baryon asymmetry according to the previous works Turok:1990zg ; Cline:1995dg ; Fromme:2006cm ; Cline:2011mm ; Enomoto:2021dkl ; Kanemura:2023juv ; Aoki:2022bkg ; Dorsch:2016nrg . The departure from thermal equilibrium is achieved by the strongly first-order EWPT. In the model, EWPT can be strongly first order by the non-decoupling effect of the additional scalar bosons, which also induces the sizable deviation in the Higgs triple coupling Kanemura:2002vm ; Kanemura:2004ch ; Kanemura:2004mg ; Braathen:2019pxr ; Braathen:2019zoh . In our benchmark scenarios, by taking $\mu_{S}=20~{}\mathrm{GeV}$, we have $R_{hhh}\equiv\Delta\lambda_{hhh}/\lambda_{hhh}\simeq 22~{}\%$, where we used the one-loop effective potential to evaluate $\Delta\lambda_{hhh}$. According to Refs. Turok:1990zg ; Cline:1995dg ; Fromme:2006cm ; Cline:2011mm ; Enomoto:2021dkl ; Kanemura:2023juv ; Aoki:2022bkg ; Dorsch:2016nrg , this would be large enough to make EWPT strongly first-order. Consequently, we expect that EWBG can work in our benchmark scenarios. Since a more detailed study needs complicated analyses using the effective potential, we will study it in our next paper to avoid making this paper too long future .
The non-decoupling effect in $R_{hhh}$ can be tested in the process of di-Higgs production at future high-energy colliders Cepeda:2019klc ; Fujii:2015jha ; Bambade:2019fyw . $R_{hhh}$ in the benchmark scenarios would be tested at the upgraded ILC with the beam energy $\sqrt{s}=1~{}\mathrm{TeV}$, where the expected sensitivity is $10~{}\%$. Strongly first-order EWPT can also be probed by the gravitational waves Grojean:2006bp ; Espinosa:2010hh ; Kakizaki:2015wua ; Hashino:2016rvx ; Hashino:2018wee at future space-based gravitational wave observatories LISA:2017pwj ; Seto:2001qf ; Corbin:2005ny . Furthermore, a sizable non-decoupling effect in EWPT can be tested by the observations of the primordial black holes Hashino:2021qoq ; Hashino:2022tcs
@@ Acknowledgements
The work of K. E. was supported in part by the National Research Foundation of Korea (Grant No. RS-2024-00352537). The work of S. K. was supported in part by the JSPS KAKENHI Grant No. 20H00160 and No. 23K17691.